The manager of a coffee shop has one type of coffee that sells for $6 per pound and another that sells for $9 per pound. The manager wishes to mix 40 pounds of the $9 coffee to get a mixture that will sell for $7 per pound. How many pounds of the $6 coffee should be used? a.40lb b.60lb c.80lb d.120lb
Found 2 solutions by richwmiller, Nate:
Answer by richwmiller(17219)

40*9+6x=7*(40+x)
x=80 lbs.

Answer by Nate(3500)

The average of the prices will equal 7.00
( 40*$9 + x*$6 ) / (40 + x) = $7
x = 80lb